

********************************************************************
UARS WEEKLY STATUS REPORT - 03 NOV 2000
********************************************************************
>>>>> 50,000th ORBIT <<<<<
The UARS has operated over nine years and completed its 50,000th orbit
today. GO UARS! (next announcement at orbit 100,000)
***********
SOLSTICE
***********
SOLSTICE is in good health, collecting solar data on all available orbital
tracking periods and stellar calibration observations whenever possible.
Our solar data is unaffected by the larger than normal pointing errors
since they are within our calibrations.
Additionally, we continue to gather enough data to meet all the SOLSTICE
science objectives during this period of reduced tape recorder operations.
However, level2 processing is halted due to deficiencies in the OA_SSPP_ATT
service. CDHF and CSC personnel released the engineering patch to this and
associated sevices and we are re-running our tests.
***********
SUSIM
***********
SUSIM continues to gather a normal set of daily solar scans scheduled
for predicted periods of UARS telemetry.
We are preparing to begin the reprocessing the SUSIM UV spectral
irradiances to data version V21 in the next few weeks.
***********
ACRIM
***********
Nothing received.
***********
PEM
***********
The PEM zenith spectrometers are activate in daylight mode.
At the present time, PEM energy deposition from its spectrometers is not
being generated. Recent data configuration changes have affected PEM software
causing multiple failures. These failures are being reviewed by the PEM team.
In order to correct for new telemetry requirements, PEM will need to update
processing code with a patch delivery. The delivery patch may include changes
which would allow the PEM software to execute on a DEC Alpha machine. GSFC is
helping PEM with some DEC Alpha particulars which have caused PEM level 3
code to fail on DEC Alpha machines (software below level 3 executes on the
DEC Alpha).
The PEM team is working on replacing the PEM:: machine with a DEC Alpha
machine to be called PEM2K::. PEM:: is now over 10 years old and its hardware
is now very difficult and costly to maintain. PEM:: is also showing signs
of slowly failing, and it is doubtful that PEM:: could be repaired if it failed.
The PEM system depends on PEM:: for instrument commanding and telemetry flow,
so it is an integral part of the PEM ground system.
Updated HEPS calibration parameters have been finalized and are ready for
delivery. Updated MEPS quality calibration files have been finalized and are
ready for delivery. Level 3 software errors, discovered by Dr. Dean Pesnell,
have been fixed; his suggestions have been incorporated with in the PEM
production code. Both updated HEPS and MEPS calibration data, along with the
Level 3 code changes, will influence the in-situ energy deposition predictions,
so new versions of PEM energy deposition Level 3 files will be generated.
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
The MPG participated the UARS Attitude Status control meeting on
1 November 2000.
For the benefit of the UARS community, here is a summary of the
status reported at that meeting: Engineering tests to simulate
onboard processing of Three-Axis Magnetometer (TAM) data have
been very encouraging. Tests with the UARS simulator suggest
that onboard processing of TAM data will maintain spacecraft
attitude with sufficient accuracy for all UARS instrument
observations and their science applications. Development is
ongoing to code and test the software changes which will be
uplinked to the UARS onboard computer. Until testing of the
flight software is completed, the spacecraft will continue to use
the Fine Sun Sensor (FSS), backed up by ground processed
magnetometer data, for attitude control.
When the solar beta angle increases to above 64 degrees the Sun
is not in the FSS field of view. UARS will encounter this
condition from 7 to 16 November. During this status meeting it
was recommended that UARS operate with a yaw offset of +10
degrees. With this 10-degree yaw offset the FSS will continue to
observe the Sun and provide updates for attitude control during
this period of high solar beta angles.
The MPG also participated in the UARS Monthly Status Review
meeting on 1 November. Daily processing of instrument Daily
Activity Plans continues with no problems. Instrument operations
schedule is complete to 3 December 2000. The MPG continues to
coordinate with the Flight Operations Team and to assist the
Project Scientist with the scheduling of instrument operations.
***********
CDHF
***********
HIGHLIGHTS
- A new UOAS (per Solstice request) is available.
- Data created in the month of August 2000 for HALOE, NMC, and UKMO has been
loaded into the DAAC Transfer Tables. DAAC has been notified.
- MLS Version 5 data is being previewed and loaded into the DAAC Transfer
Tables. Various fields in the DAAC Tables are being updated that are
needed to meet DAAC criteria. DAAC (James Johnson) will be notified
when he is able to reveal/make available to User Community.
- WINDII requested Version 11 be sent to the DAAC from start of mission to
31-Aug-1997. WINDII needs to update the UARS_QUALITY field. Currently
waiting response from WINDII that the fields have been updated. DAAC
will be notified.
UARS CDHF REHOST:
.PI validation of output products and concerns are be addressed.
.Performance criteria is being validated.
***********
FOT
***********
REPORTING PERIOD
This report covers 21 October 2000 (Orbit 49801, GMT Day 295)
through 27 October 2000 (Orbit 49905, GMT Day 301).
SPACECRAFT OPERATION
The observatory is now in Reverse Flight (Northern Hemisphere
viewing) and is performing nominally under two-battery operations.
The instrument operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
HRDI 295/0056 49801 LOW Power
ACRIM 295/2318 49815 Power OFF
SUSIM 295/2318 49815 Power OFF
ACRIM 298/2337 49860 Power ON
SUSIM 299/0020 49861 Power ON
SUSIM 299/1540 49870 Power ON
PEM 300/1219 49883 DAY ONLY OPS
PEM 300/1316 49884 AXIS1 ON
HALOE 301/0023 49891 Power ON PART 1
HALOE 301/0858 49896 Power ON PART 2
A Forward-to-Reverse Yaw Around maneuver (number 93) was
successfully executed in Orbit 49299 beginning 297/11:30:00 and
completing at 297/14:10:00 on 23 October 2000. The SOLSTICE
instrument was in day only and all other intruments were in their
off or standby mode (including HRDI and PEM). Due to the SA
position at 269 degrees, the effective night length of the
spacecraft after the maneuver was approximately 54 minutes. Post
maneuver analysis of data indicated that the end-of-night load bus
voltage was 24.92 volts and the average depths of discharge (DODs)
on Batteries 2 and 3 were approximately 21.09 and 20.97 percent
respectively. The next (reverse to forward) YAW is scheduled for
02 December 2000 (DOY 337).
The beta angle ranged between the angles of 8.9 degrees to Ð19.1
degrees for this report period. The beta angle is now increasing
toward a maximum angle of 75 degrees on 12 November 2000 (DOY 317).
Spacecraft battery 1, 2, and 3 performance monitoring continues.
Battery 1 remains OFF the charge relay. Battery 2 and Battery 3 are
currently maintaining greater than 23.8 V end-of-night (EON) load
bus voltage. Battery temperatures are stable with a temperature
delta between Battery 2 and Battery 3 of 1.6 to 1.9 degrees C.
Additional battery performance data for each battery on 20 October
2000 and 27 October 2000 are provided in Appendix D.
The clock error ranged from Ð9.0 to 5.0 msec during this report
period. Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT
for this report. A summary is listed below and the details of ALERT
and Out-Of-Limits occurrences are listed in Appendix B.
¥ OUT-OF-LIMITS = 01
¥ ALERTS = 00
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
¥ FOT = 11
¥ HRDI = 01
¥ SUSIM = 03
¥ ACRIM = 02
¥ HALOE = 02
¥ PEM = 03
UARS Anomaly Reports initiated during this report period are listed
in Appendix E and summarized below.
¥ FOT = 02
One (1) Anomaly Report remained OPEN at the end of this report
period (see Appendix E).
TRANSPONDER FREQUENCIES:
- Transponder A: 00/221 - 2287.496937
- Transponder B: 00/270 - 2287.499037
GROUND SYSTEM OPERATION
Two (2) operational support problems occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF)
from mission start through UTC Day 00/300 show the observatory data
loss to be 298 hours, 45 minutes, 15.680 seconds (an increase of 5
minutes 20.016 seconds since the last report period). This is a
0.39 percent data loss which equals a 99.61 percent data capture for
the mission.
The increase of 5 minutes 20.016 seconds occurred over three (3)
days:
- 1 minute 18.144 seconds on Day 294 due to a late AOS,
- 3 minutes 25.440 seconds on Day 295 due to a late AOS,
- 0 minutes 36.432 seconds on Day 300 due to a data hit.
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
Testing continues daily to get the TAM and FSS attitude control
solution operational. The UARS testing and training simulator
(UTTS) is being used for all testing. In order to insure the
validity of the test the FOT has been debugging some problems with
the simulator attitude models. This week two bad bits were found in
the FSS data on the simulator. The FOT is attempting to isolate the
problem.
MEETINGS
The FOT attended a meeting to discuss UARS Attitude on 27 October
2000.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for December 02, 2000 (DOY 337).
This maneuver will be a Reverse to Forward Flight maneuver (Northern
Hemisphere viewing).
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 27 October 2000
ORBIT AOS NO. SUBSYS REASON
------- ------- ----- -------- ---------
49801 295/0056 00-582 HRDI LOW Power
49808 295/1117 00-583 PEM PEM GROUP DISABLED
49815 295/2318 00-584 ACRIM Power OFF
49815 295/2318 00-585 SUSIM Power OFF
49837 297/1020 00-586 ACS FWD Flight GYRO BIAS
49837 297/1020 00-587 ACS REV Flight GYRO BIAS
49838 297/1120 00-588 ACS YAW MANEUVER - REV FLIGHT
49840 297/1452 00-589 OBC/ACS FSS TO UPDATE MODE
49851 298/0854 00-590 NBTR PLAYBACK
49854 298/1358 00-591 OBC GYRO BIAS UPDATE
49860 298/2337 00-592 ACRIM Power ON
49861 299/0020 00-593 SUSIM Power ON
49870 299/1540 00-593 SUSIM Power ON
49883 300/1219 00-594 PEM DAY ONLY OPS
49884 300/1316 00-595 PEM AXIS1 ON
49885 300/1459 00-596 OBC GYRO BIAS UPDATE
49887 300/1835 00-597 NBTR NOOPS D301
49891 301/0023 00-598 HALOE Power ON PART 1
49894 301/0521 00-597 NBTR NOOPS D301
49896 301/0858 00-599 HALOE Power ON PART 2
49900 301/1502 00-601 OBC GYRO BIAS UPDATE
49903 301/2018 00-600 NBTR NOOPS D302
________________________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 27 October 2000
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
--------- ------- ------------ --------
OBSZ1OOL UFLTR Z1 Residual Exceeded 100 Arcsec 49840
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
--------- ------------ --------
-None this report period
________________________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 27 October 2000
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
------ ----- -----------------
N/A N/A N/A
____________________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 27 Oct 2000
20 Oct 2000 - GMT Day 294 Beta = 12.8 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ----------- -------- ---------- ----- ------- ------
1 +728.0/+728.0 -3.68/-4.94 +0.0/+0.0 14.2 0.000 0.0
2 +22.4/-5.6 +4.51/+3.87 +29.2/-12.8 25.8 1.040 19.4
3 +00.0/-95.2 +3.55/+2.60 +28.4/-12.8 25.8 1.040 19.0
27 Oct 2000 - GMT Day 301 Beta = 19.1 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
---- ----------- --------- ------ ------ ------ ----
1 +728.0/+728.0 -3.05/-4.94 +0.0/+0.0 14.2 0.000 0.0
2 +56.0/-11.2 +5.79/+3.87 +30.4/-14.4 24.2 1.030 21.5
3 00.0/-123.2 +4.51/+2.92 +29.6/-15.2 24.2 1.020 21.4
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
-------- ------- --------------
no changes V/T 5
________________________________________________________________________
APPENDIX E
UARS ANOMALY REPORTS
Week ending 27 October 2000
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ------- ----- ---- ------------------
00-042 FOT 49810 295/1437 14170 Generic Late Acquisition
00-042 FOT 49810 295/1520 14170 Gen
LOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
----- -------
-None this report period
ANOMALY INVESTIGATION REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
----- ------ ----- ----- ---- ------------------
-None this report period
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
----- ------ ----- ----- ---- ------------------
00-042 FOT see list Oct 00 14170 Generic Late Acq.
SEVNTFW00/43